Maintenance and repairs log — Redsail Laser

Repair, maintenance and ventilation history for the Redsail Laser. Routine maintenance procedures and troubleshooting are on Redsail Laser — Maintenance instructions.

Entries are grouped by year, newest first. Dates containing xx or an asterisk are incomplete and have been retained as recorded.

2026

2026-10-14

Laser would not fire because the compressor had tripped on HIGH TEMP. Restarting the compressor restored air supply and laser operation. Slack

2026-10-04

Test cuts were out of square by about 1 mm over 100 mm, with X dimensions also short. No correction was confirmed. Slack

2026-09-20

Z-table movement failed again; both driver LEDs were off. Repeated restarts temporarily restored movement, but no repair was recorded. Slack

2026-09-11

Alvar reported mirror alignment and tube repositioning a few weeks earlier. Good cutting was confirmed; the exact repair date was not given. Slack

2026-08-22

A fallen LED strip blocked the beam across the middle of the bed. Repositioning it restored cutting. Slack

2026-08-07

Bullen repaired a broken lid gas-strut mounting bracket by welding in a new piece of angle iron. Slack

2026-07-18

Bullen replaced the filter. The xTool table had to be moved to reach and remove the filter holder. Slack

2026-06-01

Mirrors were realigned again after alignment shifted during locking. The tube had been raised 2 mm on 31 May; a possible first-mirror beam offset remained. Slack

2026-04-17

Coolant was below the sight glass and air was circulating in the cooling system. An emergency refill was planned. Slack

2026-04-14

An out-of-range Z coordinate blocked table movement. Jogging it back into range restored movement; direction-signal wiring remained suspect. Slack

2026-04-11

Replaced filter. Fixed connector for the Z table. Cleaned and greased the table legs.

2026-02-26

An xTool extraction fault improved after a filter change, but the Redsail red beacon still blinked when the xTools were powered on. One-at-a-time use resumed. Slack

2026-01-31

Repaired the shorted high-voltage laser power supply (Cloudray Z100). After repeated testing, the machine’s blown fuse switches were traced back to the laser power supply unit. The fault was identified as a failed filter capacitor (C1), which resulted in inadequate suppression of transient voltage spikes. This caused excessive transient voltages to pass through the circuit, ultimately leading to failure of the protective TVS diode, which shorted the supply as designed. Replaced the following components: C1 capacitor (400 V, 220 µF) and TVS diode (1.5KE440CA, supplied by ELAB). No additional component damage was detected, indicating that the protection circuitry successfully prevented further failure.

2026-01-02

Z-table movement failed and both X/Y end switches needed adjustment. No completed repair was reported. Slack

2025

2025-12-09

After no beam reached the bed, Robert removed debris from the beam opening and Bullen reseated a lens that was not lying flat. Cutting was restored that evening. Slack

2025-12-07

Lens and mirrors cleaned. Aligned mirrors. Raised the laser tube. Moved the second mirror towards the left.

Damaged cables had already been removed. Cutting remained slightly weak, and intermittent X-limit warnings were reported. Slack

2025-12-03

Smoke and damaged wiring near the laser head were reported. The laser was switched off and labelled out of service. Slack

2025-11-23

Running the Redsail and the new duct fan together triggered a low ventilation-pressure warning. Suction remained, but the cause needed further investigation; no repair was recorded in this thread. Slack

2025-11-19

Filter replaced

2025-08-31

Daniel reported squealing from the chiller on startup and a stuck Z table. The cause and repair were not established in the thread. Slack

2025-08-19

Filter replaced

Estrid reported that the filter was heavily clogged; a subsequent run showed no smoke leakage. Slack

2025-08-18

Following a small fire while cutting thick acrylic on 17 August, David cleaned the bed slats and returned the laser to use. Cutting still seemed weak; no definite fire damage was reported. Slack

2025-08-12

Behzad realigned the mirrors, checked all four corners and different Z heights, and reported clean engraving and cutting in the test run. Slack

2025-06-09

The focus adjustment nut was reported stuck. Bed-height adjustment allowed work to continue, but the thread records no repair to the nut. Slack

2025-05-30

Jonas realigned all three mirrors and cleaned the mirrors and lens. Test cuts produced narrow, precise lines; a few lens blemishes remained. Slack

2025-05-16

Zik adjusted the mirrors, including the head mirror whose locking was loose. Cutting improved but remained below normal. Slack

2025-05-11

The centre LED strip in the lid was hanging down. It was pushed back into place as a temporary measure; permanent fastening was only proposed. Slack

2025-05-07

Christian Antila cleaned the lens and adjusted the second mirror, improving poor cutting on the right side of the bed. The head mirror still needed attention. Slack

2025-05-03

Alvar replaced a damaged focusing lens after very poor cutting was reported. Lens orientation was discussed and changed afterward; no completed cutting test was recorded. Slack

2025-04-24

Alex raised the laser tube and fitted a new lens labelled 63.5 mm after an unsuitable lens had been used during testing. A quick cut worked; further mirror-height adjustment and alignment were still recommended. Slack

2025-04-05

Filter replaced

2025-03-29

Bullen repaired a homing limit switch by adding a spacer underneath it after the head failed to stop during reset. The repair was confirmed that evening. Slack

2025-03-20

Klas replaced the broken X-axis limit switch and bent its actuator to fit. Homing was tested successfully, and the laser was returned to service. Slack

2025-02-05

Pontus secured a loose cable away from the moving gantry using a printed cable-tie holder. Slack

2025-02-04

Second mirror fixture replaced. Screws that hold the boom tightned. Lens replaced. Mirrors aligned. Filter replaced.

The laser returned to use, but cutting was still slightly weaker in the upper-right corner. Slack

2025-01-29

Aron calibrated the optics and cleaned the lens during the 28–29 January troubleshooting. The laser worked again, but cutting remained weaker than normal. Slack

2025-01-28

Troubleshooting found the pulse power set to 11%; correcting it restored the pulse function. A properly timed measurement gave about 94 W directly after the laser tube. Slack

2025-01-27

The air-assist hose at the gantry connection repeatedly came loose. A worn hose or fitting was suspected; replacement was discussed but not confirmed. Slack

2025-01-14

Oscar Wahlberg realigned the mirrors after the laser became too weak to cut. Normal operation was reported restored. Slack

2025-01-09

Lens replaced, but it's not the correct type. We ran out of the normal ones. Mirrors aligned.

2025-01-06

Lid for the laser was fixed and strenghted.

2024

2024-12-15

Jonas cleaned the mirrors after the pulley and head work. The laser cut in all four corners again, although the lower-right beam remained more dispersed. Slack

2024-12-14

Replaced laser head. Greased all the grease nipples except one on the right Y-axis that is not possible to reach. X-axis left pulley assembly was fixed after it broke a few weeks earlier. New bearings for it. New axle and stop to prevent the belt from leaving the pulley was made on the lathe.

A longer X-axis grease nipple was fitted; noisy bearings and the inaccessible rear Y-axis lubrication point were noted. Slack

2024-12-03

Didrik restored the failed yellow status light with a temporary repair. Replacement of all its LEDs remained proposed. Slack

2024-11-20

Cleaned mirrors and lens. Replaced filter. Aligned mirrors. Laser power was measured to 90W just before the first mirror.

Bearings and rails were also lubricated; cutting remained uneven despite the alignment checks. Slack

2024-10-28

Edvin adjusted the head and replaced lens seals after recurring poor cutting. Operation improved temporarily, but the beam still changed with head rotation. Slack

2024-10-18

Jonas cleaned the lens, replaced its seals and adjusted the mirrors. Alignment remained incomplete, and subsequent attempts still hit the nozzle. Slack

2024-09-18

Replaced filter.

2024-08-28

Repeated circuit trips were reported, and the storage shelves were moved to allow an electrician access to the suspected wall outlet. No completed outlet repair was recorded in this post. Slack

2024-08-17

Daniel and another member adjusted the mirrors after cutting was limited to the upper-left corner. The work was confirmed the following morning. Slack

2024-07-03

Replaced filter.

2024-06-29

Cleaned mirrors and lens.

Daniel tested ventilation startup without the wood-room button and reported it working. Slack

2024-06-19

Ventilation modifications were completed after the May shutdown for fumes entering the building. A roughly 15-minute acrylic-cutting test passed without smell in the offices; the out-of-hours manual-start issue remained. Slack

2024-06-02

Daniel reported successful cleaning of the bed slats after soaking them with laundry detergent. Slack

2024-05-22

The laser was taken out of service because its fumes were spreading into the rest of the building. Ventilation modifications followed in June. Slack

2024-05-01

Replaced filter.

2024-03-13

Cleaned mirrors and lens.

2024-02-16

After a reset with the controller's rotary setting still enabled caused a Y-axis crash, Andy Yu realigned the laser and restored cutting. Slack

2024-02-14

Balázs wiped and refitted half the removable bed slats after Daniel soaked them overnight. The other half remained uncleaned. Slack

2024-02-09

Christian Hanssen printed a cap to allow larger-diameter cylinders to be mounted in the rotary attachment. Slack

2024-01-23

Daniel restored cutting acceleration after it had been set almost to zero. Cutting motion recovered, although the original acceleration values were not known exactly. Slack

2024-01-06

A former CAD computer was installed beside the laser, with LightBurn working. Slack

2024-01-03

David tightened the loose front laser-tube mount screws and retightened the rear screws after the December alignment problems. Cutting worked across the whole bed with the usual settings. Slack

2023

2023-12-26

Aron and Annica recalibrated the optics for about two hours. The laser was reported working again. Slack

2023-12-21

The laser was taken out of service after ventilation flow was lost during building ventilation work by Akademiska Hus. The red warning remained active and an out-of-order notice was posted. Slack

2023-12-08

Replaced lens. Levelled the machines frame. Adjusted squareness of x-y. Adjusted mirrors.

Mika also secured the loose first mirror pillar during the 7–8 December work. Cutting power and angle were reported correct. Slack

2023-12-06

Replaced end switch for y axis. Replaced cables to x axis motor controller, because it only went in one direction. It seems that the controller didn't get the direction signal

2023-10-18

Andy adjusted the gantry alignment and realigned the optics to reduce out-of-square cuts. The measured angular error improved from about 0.48° to 0.2°, so further correction was still needed. Slack

2023-10-10

Edvin shortened folded belt ends by a few centimetres after they were found interfering with travel. No dimensional retest was recorded in the thread. Slack

2023-09-16

Switched the power supply to a new one, the new one is Cloudray Z100

Replacement followed reports of fading output during longer jobs. The digital link to the controller remained unconnected pending a suitable terminal connector. Slack

2023-09-15

A low-ventilation-pressure alarm was traced to increased airflow through an opening in the new wood room. Temporarily blocking the opening cleared the alarm; permanent restoration of the silencer box was still planned. Slack

2023-08-30

Y-axis end stop switch was bent so the laser crashed and scraped against the steel plate in the back. Bent back the end stop switch. Crash caused mirrors to go out of alignment. Adjusted and cleaned mirrors. Laser seemed to shake when starting moves between operations so lowered start speed and acceleration for those kind of moves.

2023-08-13

The yellow status light appeared faulty; the red light worked when tested, suggesting a control-signal problem. The X-axis grease nipple was obstructed by the belt, and the power-supply replacement remained pending. Slack

2023-08-09

Filter replaced, lubricated rails, moved switch for x-axis to prevent crashes, lens replaced (it was fine, we dropped it), fixed issue with door beneathe laser not closing properly.

2023-07-19

The tube stopped firing while the rest of the machine still operated. Mischa and Mats investigated; the black activation button tested correctly, and a replacement power supply was available but not yet fitted. Slack

2023-05-23

Laser tube height adjusted, mirrors adjusted, lens replaced.

2023-05-18

Niclas corrected the lens assembly; Mischa reported aligning the first and Y-axis mirrors. Cutting resumed, but angled cuts and head alignment needed further work. Slack

2023-03-10

Didrik traced the laser failing to fire to lost compressed-air pressure after the compressor breaker tripped. Laser operation was restored; the cause of the compressor fault still needed investigation. Slack

2023-02-19

Edvin confirmed that the damaged mains connector had been repaired. The lead was left unplugged to prevent another trip over it. Slack

2023-01-30

Network connectivity was restored by moving the Ethernet connection to another wall port. The original sockets remained suspect; rerouting was proposed after the machine relocation. Slack

2022

2022-12-14

Niclas cleaned and aligned the mirrors, corrected the lens orientation and mounting, restored its missing seal and repositioned the head mirror. Ghosting was removed and MDF cuts improved, although slower-than-usual speeds were still needed. Slack

2022-12-05

Daniel carried out a brief optical realignment to improve cutting. Slack

2022-10-23

The extraction filter was replaced. Didrik soldered loose ventilation-control wires after forced extraction stopped working; operation worked on the daytime schedule, but the forced-start repair still needed an evening test. Slack

2022-10-11

Mika found loose screws on the second mirror. Daniel and Mats secured them, moved the first and second mirrors about 1 cm inward to prevent a screw hitting the side, and recalibrated the optics. Slack

2022-10-02

Head-mirror readjustment was reported. Checks did not explain recurring misalignment; poor cutting returned by 5 October. Slack

2022-09-23

Edvin and Bullen investigated an electrical fault and found a 230 V connector that may have been rubbing against a heatsink. They reported that they believed the problem was resolved. Slack

2022-09-11

Edvin repaired the yellow status light after failed LEDs were found in its homemade LED assembly. Slack

2022-09-09

Mirror mounts and the head were replaced during 3–9 September. Alignment was completed on 9 September with good cutting; the changed head required a new focus gauge. Slack

2022-08-15

Mika tightened loose screws securing the first mirror pillar near the laser tube and realigned the optics. The mount no longer wobbled, and test cutting was acceptable. Slack

2022-08-12

Erik, Holger and Erasmus aligned the mirrors and replaced the lens. Cutting was reported good again. Slack

2022-08-09

Filter replaced

2022-06-12

Didrik, Daniel and Bullen reported restoring the compressed-air pressure safety interlock, preventing the laser tube firing when compressed-air pressure is lost. Slack

2022-05-12

The extraction filter was replaced. Slack

2022-05-08

Daniel aligned the mirrors and reduced the working area to 139×89 cm to prevent head collisions. Tube output measured about 100 W, but cutting remained weak in the lower corners. Slack

2022-04-30

Joel adjusted the bed and aligned the mirrors, reporting good cutting. Poor cutting was reported again by 3 May. Slack

2022-03-30

Daniel and others replaced a lens that could not be cleaned, which appeared to resolve the cutting problem. The mirrors looked acceptable but could benefit from further adjustment. Slack

2022-02-20

Didrik installed a replacement 24 V control power supply after the old supply failed. The laser was returned to service. Slack

2022-01-19

Daniel replaced the lens. Cutting was thought close to normal, possibly still slightly weak; small compressed-air leaks remained at several points. Slack

2022-01-12

Daniel cleaned the lens and aligned the mirrors. A central mark on the lens would not clean off, and an acrylic test still required slower cutting than previously. Slack

2021

2021-12-27

Daniel replaced the extraction filter. Slack

2021-11-16

The failed extraction fan was repaired by an authorised person, restoring ventilation. Laser operation was confirmed again on 18 November. Slack

2021-08-22

Erasmus wiped the interior and bed blades and cleaned debris from the bottom. Deposits still needed more work; Bullen continued scrubbing the blades the next evening. Slack

2021-08-19

Mika cleaned up the focus-lock nut thread so it tightened without binding, and tightened the upper head section to prevent it unscrewing during adjustment. Slack

2021-08-08

Pär reported recalibrating X-axis distance in LightBurn the previous day after an error of about 3 mm over 1 m. Results improved; the cause was unknown. Slack

2021-08-04

Didrik cleaned the bed blades, their supporting rails and nearby surfaces, then readjusted the supports. About 1–2 mm height variation remained across the bed; test cuts looked acceptable. Slack

2021-05-26

Erik reported replacing both head tubes, nut, lens, O-rings and mirrors 2–3 after misalignment scorched the head. Cutting was usable; further alignment was needed. Slack

2021-05-16

Daniel updated engraving compensation offsets and cutting settings. Z-axis speed and acceleration were reduced to investigate recurring bed skew; X/Y start speeds were increased from 5 to 10 mm/s and tested. Slack

2021-05-12

Didrik repaired the yellow status light after a failed LED was found. The yellow segment was working again. Slack

2021-01-18

Daniel reinstalled the repaired CW-5200 chiller. Its flow alarm was initially unreliable but was reported working again later that day; the cause of the alarm recovery was unknown. Slack

2020

2020-12-26

Replaced the CW5200 pump and 24 V power board. Water flow appeared normal, but the chiller still sounded an alarm; troubleshooting continued. Slack

2020-12-16

Adjusted mirror 2. Mirror 3 alignment and a 2–3 mm front-to-back bed-level difference remained to be resolved. Slack

2020-12-07

Relevelled the bed; the remaining front-to-back height difference was about 1 mm. Slack

2020-11-29

Levelled the bed after a 6–7 mm side-to-side height difference. Sloping cut edges still required investigation. Slack

2020-11-14

Connected a CW3000 as temporary cooling after the CW5200 pump failed, restoring operation. Ordered a replacement CW5200 pump. Slack

2020-11-01

Controller replaced with a Ruida, and AutoLaser replaced with LightBurn. Max speed, max acceleration and engraving offsets adjusted.

2020-09-28

Bullen added a new fire extinguisher. Slack

2020-09-12

Adjusted the Y limit-switch arm after homing drove the gantry into the rear panel. MDF and acrylic cutting tests passed; engraving was not tested. Slack

2020-09-05

Reset the gantry after a collision made one drive belt skip; squareness was restored to within about 1 mm. Realigned the optics and cleaned the lens. Slack

2020-03-08

Repaired a loose Y-axis bearing after it came free; the bearing was intact but its retaining ring appeared stretched. Slack

2020-03-06

Checked the rotary-axis calibration after the motor replacement; a full revolution corresponded to 240 mm of commanded movement in the test. Slack

2019

2019-11-xx

Laser head replaced after thread damage; exact day unknown. A Cloudray crosshair alignment tool was purchased at the same time. Slack

2019-11-26

Reconnected and enclosed the rotary-axis switch, tidied cabling, disconnected unused motor drivers and secured a loose neutral connection. Work was reported successful. Slack

2019-11-16

Replaced the servo motors with Moons stepper motors and Leadshine drivers, plus belts and some pulleys. Motion was improved and quieter; rotary switching still needed rewiring. Slack

2019-11-05

Repaired the laser-tube clamp with an epoxied M4 threaded insert and shorter screws. Replaced deteriorated clamp padding with rubber/silicone to remove tube play. Slack

2019-09-02

Cleaned the lens and mirrors and replaced a burnt O-ring. Cutting speed improved by a few mm/s; the exhaust filter was also changed. Slack

2019-04-03

Aligned the mirrors and cleaned the lens. The laser worked well after the service. Slack

2019-02-18

During 17–18 February, levelled the bed, cleaned and aligned the mirrors, and tightened their locknuts. Cuts became consistent across the work area. Slack

2018

2018-06-29

A replacement lens was reported fitted; the focus gauge no longer matched and needed checking. Slack

2018-06-08

Refitted loose X-rail screws and tightened both axis belts to reduce backlash. Distorted circles and curved diagonal cuts remained; motor recalibration was planned. Slack

2018-05-24

Adjusted the mirror in the laser head, improving the result. Slack

2018-05-09

Filter replaced

2017

2017-12-19

An exhaust test found the ready light stayed green with the outlet blocked. The PLC filter-pressure monitoring code was not yet implemented. Slack

2017-12-18

Replaced the exhaust filter. Slack

2017-07-10

Filter replaced

2017-03-30

Filter replaced

2016

2016-07-01

Call with Indukta: they will come and install a new tube towards the end of July; to be followed up.

2016-06-30

Laser cutter introduction course; everything worked well.

2016-06-14

Laser cutter introduction course; everything worked well.

2016-06-08

Mikael and Erik cleaned the mirrors and lenses. Concluded that we probably need a rolling schedule so they are cleaned at least once a month, preferably more often. One of the mirrors had been damaged.

2016-05-24

Call with Indukta.

2016-05-11

Mikael connected both hour-meter modules and changed the PLC software to disarm the laser, including air assist, when the lid is opened. Christoffer wrote a checklist for basic laser operation.

2016-05-10

Mounted the Omron hour meter on top by the control panel. /Erik

2016-05-08

Erik received two Omron H7ET hour meters, and designed and 3D printed a frame and case for them.

2016-05-07

Didrik redesigned the circuit for switching between the linear and rotary axis, now with a microcontroller instead of logic chips. The interference problem seems solved.

2016-05-06

Erik test-cut things for Stockholm Mini Maker Faire and had problems with the rotary/linear axis switching circuit activating by itself.

2016-05-05

Didrik built a circuit for switching between the rotary and linear axis with a button on the side of the laser.

2016-04-28

Niclas did test cutting and bug testing; no serious faults found.

2016-04-27

Mikael tidied the cabling and finished the PLC programming, which now needs bug testing. Erik fitted the missing magnetic switch on the lower front doors.

2016-04-24

Mikael and Erik worked on PLC programming, tidying cabling etc.

2016-04-20

Several people continued tidying up and connecting cables and hoses; Mikael worked on the PLC programming.

2016-04-14

Phone call with Anders at Indukta: reported the measured laser power and got OK for them to come and replace our tube on site.

2016-04-13

Measured the laser power with our new laser power meter. Result: 110 W directly after the tube, 100 W just before the head and 50 W after the head. The refraction fault clearly causes large losses in the last stage.

2016-04-07

Phone call with Anders at Indukta; status unchanged.

2016-03-31

Phone call with Anders at Indukta. He had not heard back from Redsail, but said he finds it hard to see anything other than a faulty tube. Asked us to try replacing the first mirror to rule it out. /Erik C

2016-03-24

Phone call with Anders at Indukta; discussed the photos. He will contact Redsail's technicians for an assessment. /Erik C

2016-03-23

Emailed Indukta photos of the strange beam pattern from the laser. /Erik C

2016-03-23

The PLC is mostly programmed, with a sensor or two left to connect. After fitting covers over the mirrors to blow compressed air across them, the laser was no longer properly aligned. While realigning, we found the cover on the last mirror melting in a strange ring pattern about 15 mm outside the beam. This is probably the cause of the "ghost lines" seen earlier when the laser was not perfectly aligned.

2016-03-16

Mecki and Theeo ran compressed air to the mirrors. Motorised dampers connected to divert the ventilation air from the laser to the workshop room when the laser is off.

2016-03-11

Theeo bought a power supply for the new LED lighting.

2016-03-09

Theeo bought LED lighting and fitted it that evening with help from Mecki and others.

2016-02-27

PLC, contactors, switch etc. fitted by Christian and Patrik; programming started. Pressure sensor fitted on the ventilation duct. Erik fitted measuring points and hoses on the filter box and routed them into the laser.

2016-02-23

Bought compressed-air couplings and hose, to change to a quick coupling at the back of the laser and proper 8 mm hose inside.

2016-02-15

Christian fitted the beacon and DIN rails and started the PLC wiring.

2016-02-10

Mikael L got an aluminium frame to mount the filter cloth in; it is being adapted and will get steel mesh on one side.

2016-01-30

Theeo welded and painted a filter box (red and black to match the laser).

2016-01-27

Adriana bought steel mesh to hold up the filter cloth in the ventilation filter.

2016-01-20

Ducts run down from the ceiling ventilation duct; spigot and 200 mm duct fan fitted.

2016-01-13

Timer in the space mounted on the board by the notice board.

2016-01-10

All controls in the attic installed; timer wired down to the space.

2016-01-09

The ventilation company finished the ductwork.

2016-01-07

Mikael L built the ventilation controls.

2016-01-06

Test filter box in OSB built by Mikael, Erik and others to test area/pressure drop and work out its placement.

2016-01-04

Västberga Ventilation started running the new ducts.

2016-01-03

Fredrik dropped off the 315 mm duct fan we will use to pull air through the ducts.

2015

2015-12-28

Västberga Ventilation reported that their technician is ill, so the work is postponed until after New Year. /Erik C

2015-12-21

Accepted the quote for rerouting the ventilation; verbally told they can probably start between Christmas and New Year.

2015-12-19

Board decision to go with ventilation rather than a filter unit.

2015-12-16

Anders at Indukta came and aligned the machine to fix the problem with "ghost lines" we had been having. He also took the filter unit back.

2015-12-16

Received a quote from Västberga Ventilation for rerouting the ventilation ducts.

2015-12-12

Announced in the Facebook group and on the wiki that the laser cutter is out of service until further notice; started dismantling the filter unit.

2015-12-11

Fourth laser course. The last filters in the filter unit had to be fitted just before, to be able to run at all without filling the space with smoke.

2015-12-10

Anders at Indukta agreed to take back the filter unit and refund the money. /Erik C

2015-12-10

Smoke from the laser set off the fire alarm in the space again, this time while testing the ventilation. This also cost about 4000 SEK. :-(

2015-12-09

Smoke from the laser set off the fire alarm in the space; the call-out fee was about 4000 SEK. :-(

2015-11-29

Received a quote for a Purex 2000i filter unit from Alvitiq: even more expensive; still looking for cheaper options.

2015-11-27

Received a quote for a Bofa AD1500iQ filter unit from Vodex.co.uk: expensive.

2015-11-15

The filter situation is untenable; filed a complaint with Indukta and asked for a refund. /Erik C

2015-11-11

Complaint from THINGS that the building smells of smoke. /Erik C

2015-11-09

Third laser course; smoky despite newly replaced filters.

2015-11-02

Second laser course; a lot of smoke.

2015-10-30

First laser course; clearly smoky.

2015-10-28

Mikael L measured the flow through the filter unit: under 300 m³/h rather than the specified 1500 m³/h, which explains the problems. /Erik C

2015-10-20

Spent an evening aligning the mirrors. Emailed Indukta to complain about "ghost lines" beside cuts and uneven engraving.

2015-10-1*

We have cut quite a lot with the laser and realise it gets fairly smoky in the space, even though the filter (which is incredibly loud) should handle it.

2015-10-08

The laser cutter was delivered, lifted into the space, and had its mirrors aligned by Anders at Indukta.

2015-09-15

Told that delivery is slightly delayed.

2015-08-25

Delivery ETA given as 21 September.

2015-07-02

The laser cutter is ordered: a Redsail CM1490 from Indukta Technologies AB.

2015-06-30

Stretch goal reached: buying a honeycomb table and a rotary attachment as well.

2015-06-23

Crowdfunding goal reached.

2015-06-21

Crowdfunding started

2015-06-xx

Board meeting decided that the association contributes up to 100,000 SEK to buy a laser cutter and filter unit, and that the rest (just over 38,000 SEK) is crowdfunded.

2015-05-21

Received a quote for a filter unit, to run the laser indoors without separate ventilation.

2015-05-xx

Board meeting decided not to import the machine ourselves, so that there is a Swedish company to deal with if something goes wrong.

2015-04-30

Received a quote for a laser cutter from Anders at Indukta Technologies AB.

2015-01-27

First quote for a laser cutter, directly from Redsail, for importing it ourselves.

Adding a maintenance entry

Add a dated entry under the appropriate year, newest first. Include the fault or check, work carried out, parts used, who performed the work and the result. Add useful photographs and links. State whether work is proposed, in progress or completed, and whether the machine is available for use or remains out of service. Record any operating test and its outcome.

For filter changes, record the date here and on the label by the filter.